OpenStudy (anonymous):

To verify that x = 5 is a solution for the equation 8x – 13 = 2x + 17, what is a possible first step? A. Multiply 8 by 5. B. Subtract 13 from both sides. C. Replace x with 5. D. Add 17 to both sides.

OpenStudy (sweetburger):

c

OpenStudy (anonymous):

c replace x with 5

OpenStudy (anonymous):

b and d dont make sense because you would have to add 13 and subtract 17 plus thats not the first step and a makes sense but c makes the most sense

OpenStudy (sweetburger):

torch u dont have to add or subtract anything...

OpenStudy (anonymous):

c replace x with 5 and if each side equal the other then the solution is correct that x=5 and if one side doesn't equal the other the solution is wrong

OpenStudy (anonymous):

if you replace 5 for x you rewrite the equation as 8(5)-13=2(5)+17 and then you solve...

OpenStudy (anonymous):

8(5)-13=2(5)+17 40-13=10+17 27=27 the sides of the equation equal each other so the equation is correct x=5 :)

OpenStudy (anonymous):

27=27 see correct because 8 x 5 = 40-13 = 27 then 2 x 5 = 10 + 17 + 27